Timeline


and

06/11/18:

10:39 Ticket #321 (Address Dead Links on the Website) created by jacob
There are several dead links on the website, some of which point to …

06/08/18:

12:57 Ticket #278 (Need https://vso1.nascom.nasa.gov/docs/wiki to redirect to ...) closed by jacob
fixed

06/07/18:

14:25 Ticket #320 (Add Hi-C II data to MSFC data provider) created by alisdair
The Hi-C II data should be available for inclusion by the end of the year. …
14:15 Ticket #319 (Add CLASP data to VSO) created by alisdair
Emailed Amy to find out about the status of the CLASP data. […]
12:56 Ticket #13 (update MLSO data provider) closed by ed
fixed

06/06/18:

14:10 Ticket #318 (SMM / CP for years 1981 - 1989 seem to be missing from HAO DB) created by ed
1980 has data: mysql> select INSTRUMENT, count(*) from tbl_1980 group by …
13:36 Ticket #317 (Added gurman to list of trac users) closed by alisdair
fixed
13:36 Ticket #317 (Added gurman to list of trac users) created by alisdair
Created an account for Joe G. to directly log tickets.
10:27 Ticket #316 (The VSO feedback page does not work) created by alisdair
The VSO feedback page does not work. Filling out the fields and hitting …

06/05/18:

16:31 Ticket #315 (KSO data not searchable from IDL commandline.) created by alisdair
A search for KSO data works with the web interface but for some reason is …

06/01/18:

11:50 Ticket #248 (sohocat -- indexing) closed by jacob
fixed: Moved the statement that indexes id_instrume before deletes in observation.sql. Ran an update and it works fine. This was probably just a performance improvement for the soho updating script.

05/31/18:

13:51 Ticket #314 (505 errors on _every_ attemopt to download AIA synoptic (1024 x 1024) FITS ...) closed by jennifer
fixed
12:48 Ticket #314 (505 errors on _every_ attemopt to download AIA synoptic (1024 x 1024) FITS ...) created by alisdair
Joe G reports "Now getting 505 errors on _every_ attempt to download AIA …[…]
11:58 Ticket #251 (ssc_tarball error messages) closed by jacob
wontfix: I'm going to go ahead and close this guy as won't fix, after some fairly-comprehensive testing it seems like the taring code is working just fine for ssc. I can't determine why it's important for the error messages to exceed 1 kB: I'm sure there's a perfectly valid reason but it's more risky for me to meddle with the error handling code without an understanding of what the problem is than just to leave it in a presumably-functional state, especially when it comes to manual memory allocation which can quickly become troublesome.
10:10 Ticket #250 (ssc -- check recovery dates for STEREO-A) closed by jacob
fixed: Changed _update_stereo.sql so that '7' data does not show up in vso results after the end of the solar conjunction, 2015-11-09.
08:13 Ticket #290 (Weirdness w/ sot update.sh:) closed by jacob
worksforme: SOT updates normally, data comes back in queries. The database is monitored by my scripts if a problem does arise in the future.

05/30/18:

12:56 Ticket #313 (vso_info help points to Joe H. as the point of contact) created by alisdair
The vso_info routine points to Joe H as the point of contact. This should …[…]
12:53 Ticket #312 (/instruments keyword not accepted by vso_info.) created by alisdair
Despite what it says in the help (see below), the IDL command vso_info …
10:44 Ticket #311 (Establish sun drawing history) created by niles
Contact NOAA, try to establish the history of the sun drawings at …

05/29/18:

09:04 Ticket #220 (add Wendelstein Observatory data) closed by alisdair
duplicate: Closing out this ticket. It has been superseded by ticket 304.

05/25/18:

16:07 Ticket #310 (MDI data at SDAC) created by alisdair
Is the version of LOS_magnetic_field MDI data being served at SDAC the …
15:47 Ticket #309 (Attempting to download MDI data from Stanford is not working) created by alisdair
[…] also in the IDL interface […]

05/24/18:

10:03 Ticket #293 (Downloading HMI data fails when no attribute vso.Physobs is provided) closed by alisdair
fixed: Jacob's PR was accepted with modification. See https://github.com/sunpy/sunpy/pull/2621 for details.
10:02 Ticket #293 (Downloading HMI data fails when no attribute vso.Physobs is provided) reopened by alisdair

05/23/18:

09:02 Ticket #308 (Web log into database) created by niles
Just something I'm intrigued by - it may be worth ingesting the web log of …
07:19 Ticket #302 (NGDC data provider issues) closed by niles
fixed: This ticket snowballed into issues that are now better covered by tickets 305, 306 and 307. I'm resolving it because the basic issue of the many FTS files in CVS under vso/DataProviders/NGDC/indexes has been addressed. It has been established that the files are not useful and they have been removed, although I did retain 10 of them (out of the ~7K we had in CVS).

05/22/18:

14:46 Ticket #307 (SDAC GOES data) created by niles
The SDAC has GOES flare data online at : …
14:41 Ticket #306 (NCEI archival drawings) created by niles
NCEI has PDF files of sun drawings online under : …
14:35 Ticket #305 (NCEI data provider for GOES SXI data) created by niles
There are GOES SXI data for goes at NCEI, see …

05/17/18:

15:17 Ticket #304 (Serve out NOAA's sun drawings (including Wendelstein)) created by niles
NOAA has PDF files of drawings of the sun at : …
10:16 Ticket #301 (Website Search form page tweaks) closed by jacob
fixed: Patched the html so that the timing input boxes are fixed to the viewport on the right side of the page. Patched the javascript functions which operate the all day & all month buttons so that they only consider the timing data in the start row.
08:36 Ticket #196 (Annoying header on scrolling long list of records found. Header should be ...) closed by jacob
duplicate: Closing this ticket, I think the main issue is solved (the results list no longer looks like it is continually jumping back up to the top as the user scrolls). I will consider further similar changes as I move forward with #291. In its current state, the webcode is embedded in perl scripts and scattered across the filesystem, making these small aesthetic updates more time-consuming than they should be.

05/16/18:

17:16 Ticket #193 (NGDC issues.) closed by alisdair
duplicate: See ticket 302 - Niles opened a new ticked for this problem so closing this one.
16:07 Ticket #303 (HANET data provider appears to be broken) created by alisdair
Does not appear to be any response to any H-alpha net query. Seems the …
14:09 Ticket #302 (NGDC data provider issues) created by niles
Basically, there are oddities with the NGDC data provider that I need to …
08:34 Ticket #301 (Website Search form page tweaks) created by jacob
Per Alisdair's excellent suggestion, I'm going to move the date selection …

05/15/18:

11:09 Ticket #300 (Updated Trac Timeline to show updates to tickets.) closed by alisdair
fixed
11:09 Ticket #300 (Updated Trac Timeline to show updates to tickets.) created by alisdair
Edited trac.ini to set […] This enables ticket changes to be displayed …

05/14/18:

11:18 Ticket #299 (Updated distribution of ticket notifications to include Frank / Joe G / ...) closed by alisdair
fixed
11:18 Ticket #299 (Updated distribution of ticket notifications to include Frank / Joe G / ...) created by alisdair
Edited trac.ini to change the following fields. […] Tickets changes …
Note: See TracTimeline for information about the timeline view.